what are the steps to get the solution to this quadratic equation(Can you explain it to me?):
x^2 + 4x +3 =0

Answer by scott8148(6628)

factoring the expression is sort of like reverse FOILing

factors of the "c" term, 3, (±1 , ±3) that "combine" to give the "b" term, 4 ___ like 1 and 3

(x + 3)(x + 1) = 0

if the product of terms is zero, then any of the terms may equal zero

x + 3 = 0 ___ x = -3

x + 1 = 0 ___ x = -1